# Progressive Multi-Source Domain Adaptation for Personalized Facial Expression Recognition

### v1

@@ Abstract
Personalized facial expression recognition (FER) involves adapting a machine learning model using samples from labeled sources and unlabeled target domains. Given the challenges of recognizing subtle expressions with considerable interpersonal variability, state-of-the-art unsupervised domain adaptation (UDA) methods focus on the multi-source UDA (MSDA) setting, where each domain corresponds to a specific subject, and improve model accuracy and robustness. However, when adapting to a specific target, the diverse nature of multiple source domains translates to a large shift between source and target data. State-of-the-art MSDA methods for FER address this domain shift by considering all the sources to adapt to the target representations. Nevertheless, adapting to a target subject presents significant challenges due to large distributional differences between source and target domains, often resulting in negative transfer. In addition, integrating all sources simultaneously increases computational costs and causes misalignment with the target. To address these issues, we propose a progressive MSDA approach that gradually introduces information from subjects (source domains) based on their similarity to the target subject. This will ensure that only the most relevant sources from the target are selected, which helps avoid the negative transfer caused by dissimilar sources. During adaptation, the source domains are introduced in a curriculum manner. We first exploit the closest sources to reduce the distribution shift with the target and then move towards the furthest while only considering the most relevant sources based on the predetermined threshold. Furthermore, to mitigate catastrophic forgetting caused by the incremental introduction of source subjects, we implemented a density-based memory mechanism that preserves the most relevant historical source samples for adaptation. Our extensive experiments 111 https://github.com/osamazeeshan/P-MSDA show the effectiveness of our proposed method on challenging pain FER datasets: Biovid and UNBC-McMaster. Further, performance is evaluated on a cross-dataset setting (UNBC-McMaster $\rightarrow$ BioVid), showing the importance of gradually adapting to source subjects.
@@ I Introduction
Personalized FER has been extensively studied in the literature, primarily through supervised learning approaches and fine-tuning techniques [6, 7, 8] to capture individual-specific nuances. These approaches mostly rely on fully or weakly labeled data to adapt and create a personalized model for each subject. Fine-tuning a model using fully labeled data requires a costly annotation of samples and this annotation may be ambiguous due to variability among the annotators [9, 10]. Unsupervised domain adaptation (UDA) [11, 12] is a promising alternative for leveraging unlabeled data in FER. Nevertheless, SOTA UDA techniques treat datasets as domains containing samples from mixed subjects across source and target domains, which limits the model’s ability to perform fine-grained adaptation [13, 14, 15]. Defining each subject as a domain can address this issue yet blending source data to perform UDA restricts its capacity to handle variations and diversity within the target domain. To address this challenge, multi-source (unsupervised) domain adaptation (MSDA) [16, 17, 18] has gained significant popularity as it incorporates information from multiple source domains to enhance model resilience to different target variations. In [18], the authors introduced a subject-based domain adaptation method, where each domain consists of a distinct subject, thus multiple sources and a single target subject. The target domain consists of samples belonging to a single person, where the data are captured in a stationary environment and incorporate less effective diversity. This approach differs from the traditional MSDA methods that adapt to a target domain of mixed individuals. The subject-based method focused on adapting to a single individual enables more precise and targeted adaptation strategies to create a personalized FER model.
Figure 1: Comparison between the subject-based MSDA approach with our proposed progressive MSDA method. In MSDA, the alignment of domains is a crucial step that can be accomplished through Discrepancy-based, Self-supervised, Contrastive-learning, or Adversarial-based approaches. (a) In subject-based MSDA, the model takes advantage of all the source domains during the target adaptation process and performs alignment using the discrepancy-based approach. (b) Our progressive MSDA method consists of two steps. First, we rank and gradually adapt source domains (subjects) based on their similarity to the target domain, optimizing the transfer process through sequential adaptation. Second, to address catastrophic forgetting, we construct a replay memory (domain) that retains key samples from previously adapted source domains, which are re-accessed after each source-target adaptation to maintain knowledge consistency. The discrepancy-based approach is applied to align the source and target domains.
@@
In this paper, we present a new paradigm of MSDA in facial expression recognition that progressively adapts to an unlabeled target subject by exploiting the representation of the most relevant source subjects. We hypothesize that domain-level adaptation may be suboptimal in personalized FER, as the distribution of each facial expression is highly unique for each subject. To consider the most relevant sources for the unlabeled target subject, we proposed a progressive multi-source domain adaptation (P-MSDA) leveraging the closest sources for gradual adaptation, allowing a model to learn from individual source representation. Motivated by the concept of ”start small” from curriculum learning (CL) [23], we first exploit the source subjects that are closer to the target (in the feature space) and gradually introduce new subjects, while taking advantage of SPL to recalibrate and select the most pertinent sources. We compute the similarity matrix of target and source subjects while applying a predetermined threshold. This will ensure focusing only on the most relevant source subjects, as illustrated in Fig. 1(b). To train such a model, a naive approach begins with the easiest source subject and gradually moves towards the hardest subjects while storing all the previously seen subjects. Another option is to train a model by ignoring the previously learned subjects and only training with the newly added source subject. The former approach requires significant computation power and memory consumption, while the latter setting will lean towards the problem of catastrophic forgetting [27, 28]. To avoid using every visited subject and eliminate the problem of forgetting, we follow the DIL replay-based [29] strategy, where we create a replay dictionary that preserves a small set of previously adapted samples based on the closest source points from the target clusters and only incorporate those with the newly added subject. We keep updating the replay dictionary as the training progresses. Our method aims to strike a balance between leveraging diverse source information and maintaining focus on the target domain’s specific characteristics.
@@
The main contributions of this paper are summarized as follows. (1) A novel progressive learning framework for MSDA personalized FER that exploits the closest source subjects for target adaptation. Additionally, we present an effective training strategy based on DIL that starts with the closest source and gradually introduces a new subject while limiting the number of source subjects to the top-N pertinent sources. This approach minimizes the risk of data corruption from mixed source domains and reduces computational complexity. (2) Inspired by the incremental learning replay technique, we present a density-based sample selection mechanism that retains the most relevant samples from previously visited source subjects, avoiding catastrophic forgetting. (3) The performance of the proposed method is extensively evaluated on diverse groups of multiple source and target subjects on challenging BioVid and UNBC-McMaster pain estimation datasets. We also show the efficacy of our method by evaluating cross-dataset, specifically UNBC-McMaster (source) to BioVid (target). Further, we present a comprehensive analysis of the selection of previous samples for better target adaptation.
@@ II-B Domain Adaptation
Domain adaptation (DA) methods typically leverage labeled data from a source domain and unlabeled data from a target domain. The DA methods can be categorized into three main types: discrepancy-based, reconstruction-based, and adversarial-based. Discrepancy-based methods  [15, 13] aim to minimize domain shift by reducing discriminative features between domains. Reconstruction-based approaches [30, 31] focus on encoding representations of both source and target domains, often utilizing techniques such as deep reconstruction classification models and transfer learning-based auto-encoders. Adversarial-based approaches [32, 33, 34], employ adversarial techniques to adapt domain representations, such as adversarial graph representation adaptation and augmentation techniques with AC-GAN. Each category offers unique strategies to address the challenge of domain adaptation, contributing to the advancement of visual deep learning models in diverse operational contexts. Although DA-based approaches have provided a robust framework for adapting to an operational target domain. However, the reliance on a single source domain for target adaptation limits the models’ exposure to diverse data distributions. This constraint often limits its ability to have an effective target adaptation. To address this limitation, our previous work on subject-based MSDA [18] introduced a novel technique to incorporate multiple source domains (subjects) that demonstrated significant performance improvements and exhibited diversity in source domains.
@@ II-C Multi-Source Domain Adaptation
Multi-source domain adaptation (MSDA) enhances target model accuracy and robustness by leveraging multiple labeled source datasets, improving the generalization in the target domain. MSDA methods have been widely adopted in many image classification tasks [19, 35, 36, 16], these techniques can be further classified to mitigate domain shift challenges, adversarial [16, 36], self-supervised [37, 38], and discrepancy-based approaches [19]. Although we have good performance on standard benchmark datasets [35], there is still a gap in how to utilize these techniques when increasing the number of source domains in the recognition of facial expressions. Recently, Zeeshan et al. [18] proposed an MSDA method for subject-based FER, which provides a framework to tackle many source domains, i.e., up to 77, while successfully adapting to the target domain. Nevertheless, this method adapts to the target domain by leveraging multiple source subjects. While the method suggests selecting only the top-k sources, it still relies on aggregating all chosen source domains and performing joint training with the target subject. This strategy, although effective, may introduce potential noise from less relevant source domains. In contrast, our proposed strategy ranked the source domains (subjects) based on their similarity of data distribution with the target domain. Subsequently, it gradually incorporates additional subjects while continuously monitoring and optimizing performance on the target subject. This incremental integration will potentially mitigate the impact of less relevant source domains.
@@ II-E Incremental Learning
Incremental learning involves sequentially acquiring knowledge from a series of datasets without retaining access to previously learned data. It has garnered significant attention in recent years. This framework can be divided into three primary scenarios [43]: task-incremental learning [44], class-incremental learning [45], and domain-incremental learning [25]. In this paper, we focused on domain-IL for exploiting the source subjects in target adaptation. Domain incremental learning (DIL) refers to sequentially adapting to new domains without forgetting the prior information. Typically, DIL-based techniques were applied to target domains such as image classification [25], autonomous driving to detect objects [46], and semantic segmentation [47], where it is necessary to preserve the knowledge gained in the previous domain. To eliminate catastrophic forgetting, these techniques were categorized into three groups. Parameter isolation [48] stores a network or useful model parameters from previous data that is utilized with a new domain. The regularization-based strategy updates the loss by introducing regularization terms using techniques such as knowledge distillation [49, 50], which gives weight to pertinent features. Replay-based approaches [29] store some previous data by either storing them directly [49] or generating them using GAN-based techniques [28, 51]. In addition, incremental learning approaches are also applied in domain adaptation when multiple target domains exist. Wei et al. [52] proposed a distillation function to acquire previously learned knowledge. Similarly, Kiran et al. [53] adapt to multiple target domains incrementally by training a domain transfer module that generates the pseudo-images of previous targets. However, DIL-based domain adaptation methods focus on increasing target domains while retaining knowledge of the past domains, and typically overlook the domain shift problem. In contrast, the proposed technique adapts to a single unlabeled target and focuses on incrementally introducing the source domains while preserving only the most relevant samples by mitigating discrepancies between the domains to improve target adaptability.
@@ II-F Self-paced Curriculum Learning
Curriculum learning (CL) [23] and self-paced learning (SPL) [24] have been widely adopted in many machine learning and computer vision tasks. Both learning paradigms follow a structured approach, starting with simpler tasks and progressively transitioning to more complex ones. Specifically, CL leverages prior knowledge from multiple source domains to guide the learning process while SPL dynamically adjusts to the model learning pace during training. Choi et al. [54] proposed a CL-based technique for UDA. It applies CL to introduce the target pseudo-labels based on the density clustering method; it picks the subset with high density first and moves towards the lower-density samples later. Wang et al. [55] introduced an SPL-based approach that considers easy samples from the target domain based on the agreement on the predictions from the source classifiers. Initially, these classifiers are trained on the individual source domains, and then a single classifier, including all the domains, is trained. If both classifiers agree, those samples will be included first in the training process. Jiang et al. [56], propose self-paced curriculum learning (SPCL), a unified framework that integrates prior knowledge with adaptive learning progress. Formulated as an optimization problem, SPCL balances the strengths of both approaches by incorporating fixed prior knowledge and dynamically refining the curriculum throughout training. Yang et al. [57] proposed a CL-based MSDA method that relies on domain discriminator loss for the selection. The samples with a similar distribution as the target are more challenging for the discriminator to separate and are selected first. However, these approaches do not consider how the source samples were introduced; instead, they focus on presenting the most confident target samples in the training protocol. Furthermore, Yang et al. [57] technique utilizes all the samples in the source domain, which is relatively ineffective in subject-based FER scenarios as they do not consider the domain negative impact that drifts the model for adapting to a particular target subject. Hence, novel approaches are required to handle this challenge effectively.
Figure 2: Overview of our proposed progressive MSDA method for the adaptation to the target subject. Source Selection Phase: We estimate the similarity matrix between every source and target embedding, followed by ranking the sources from most to least similar subjects. Progressive Domain Adaptation Phase: Ranked sources are progressively incorporated through iterative training steps (Train Step-1, Train Step-2, …, Train Step-n). At each step, a new source subject is introduced and aligned with the target by calculating discrepancy and supervised losses. The Augmented Confident Pseudo-label (ACPL) technique from [18] generates reliable pseudo-labels for the target. Finally, we create a replay dictionary using a density-based selection to preserve previously visited relevant source samples
@@ III-A Preliminaries
In MSDA, given a set of labeled source domains $\mathcal{S}={\{\mathbf{S}_{1},\mathbf{S}_{2},\ldots,\mathbf{S}_{a},\ldots,\mathbf{S}_{D}\}}$ and a single unlabeled target domain $\mathbf{T}$, where $a={\{1,2,\ldots,D}\}$ is the number of source domains. To preserve relevant samples from previous source domains, we define a replay domain $\mathbf{R}$. We define a source domain as $\mathbf{S}_{a}={\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})}\}_{i=1}^{N^{\text{s}}}$, where $N^{\text{s}}$ represents the number of samples within each source domain $a$. Assuming that $\mathbf{x}_{i}^{\text{s}}$ represents an input embedding of a source domain sample $i$ produced by an encoder $F$, the labeled source domain input space is defined as $\mathbf{x}_{i}^{\text{s}}$, and their respective labels as $y_{i}^{\text{s}}$. We define a single unlabeled target domain as $\mathbf{T}={\{\mathbf{x}_{i}^{\text{t}}\}}_{i=1}^{N^{\text{t}}}$, where $N^{\text{t}}$ denotes the number samples within the target domain. A replay domain is defined as $\mathbf{R}={\{(\mathbf{x}_{i}^{\text{r}},y_{i}^{\text{r}})}\}_{i=1}^{N^{\text{r}}}$, where $N^{\text{r}}$ represents the number of relevant source domains from $\mathcal{S}$, which keeps on updating after adapting the model $\Theta$ to the target domain $\mathbf{T}$. The model $\Theta$ integrates the representation learning module $F$ with the classifier $C$. The objective is to leverage the information of source domains by gradually introducing domains from $\mathcal{S}$ to improve the performance of a model $\Theta$ on the target domain $\mathbf{T}$. Therefore, at each training step, the model $\Theta$ aims to learn from a new source domain from $\mathcal{S}$ and $\mathbf{R}$, which retain the most representative information from the previous source domains.

### v2

@@ Abstract
Personalized facial expression recognition (FER) involves adapting a machine learning model using samples from labeled sources and unlabeled target domains. Given the challenges of recognizing subtle expressions with considerable interpersonal variability, state-of-the-art unsupervised domain adaptation (UDA) methods focus on the multi-source UDA (MSDA) setting, where each domain corresponds to a specific subject, and improve model accuracy and robustness. However, when adapting to a specific target, the diverse nature of multiple source domains translates to a large shift between source and target data. State-of-the-art MSDA methods for FER address this domain shift by considering all the sources to adapt to the target representations. Nevertheless, adapting to a target subject presents significant challenges due to large distributional differences between source and target domains, often resulting in negative transfer. In addition, integrating all sources simultaneously increases computational costs and causes misalignment with the target. To address these issues, we propose a progressive MSDA approach that gradually introduces information from subjects (source domains) based on their similarity to the target subject. This will ensure that only the most relevant sources from the target are selected, which helps avoid the negative transfer caused by dissimilar sources. During adaptation, the source domains are introduced in a curriculum manner. We first exploit the closest sources to reduce the distribution shift with the target and then move towards the furthest while only considering the most relevant sources based on the predetermined threshold. Furthermore, to mitigate catastrophic forgetting caused by the incremental introduction of source subjects, we implemented a density-based memory mechanism that preserves the most relevant historical source samples for adaptation. Our extensive experiments  [^1] https://github.com/osamazeeshan/P-MSDA show the effectiveness of our proposed method on challenging FER datasets: Biovid, UNBC-McMaster, Aff-Wild2, and BAH. Further, performance is evaluated on a cross-dataset setting (UNBC-McMaster $\rightarrow$ BioVid), showing the importance of gradually adapting to source subjects.
@@ I Introduction
Personalized FER has been extensively studied in the literature, primarily through supervised learning approaches and fine-tuning techniques [6, 7, 8] to capture individual-specific nuances. These approaches mostly rely on fully or weakly labeled data to adapt and create a personalized model for each subject. Fine-tuning a model using fully labeled data requires a costly annotation of samples, and this annotation may be ambiguous due to variability among the annotators [9, 10]. Unsupervised domain adaptation (UDA) [11, 12] is a promising alternative for leveraging unlabeled data in FER. Nevertheless, SOTA UDA techniques treat datasets as domains containing samples from mixed subjects across source and target domains, which limits the model’s ability to perform fine-grained adaptation [13, 14, 15]. Defining each subject as a domain can address this issue yet blending source data to perform UDA restricts its capacity to handle variations and diversity within the target domain. To address this challenge, multi-source (unsupervised) domain adaptation (MSDA) [16, 17, 18] has gained significant popularity as it incorporates information from multiple source domains to enhance model resilience to different target variations. In [18], the authors introduced a subject-based domain adaptation method, where each domain consists of a distinct subject, thus multiple sources and a single target subject. The target domain consists of samples belonging to a single person, where the data are captured in a stationary environment and incorporate less effective diversity. This approach differs from the traditional MSDA methods that adapt to a target domain of mixed individuals. The subject-based method focused on adapting to a single individual enables more precise and targeted adaptation strategies to create a personalized FER model.
Figure 1: Comparison between subject-based MSDA and our proposed progressive MSDA. (a) Subject-based MSDA aligns all source domains simultaneously with the target using a Discrepancy-based, Self-supervised, Contrastive-learning, or Adversarial-based approaches. (b) Our Progressive MSDA first rank and gradually adapts source domains (subjects) based on their similarity to the target domain, optimizing the transfer process through sequential adaptation. Secondly, we construct a replay memory (domain) that retains key samples from previously adapted source domains, which are re-accessed after each adaptation. The discrepancy-based approach is applied to align the source and target domains.
@@
In this paper, we introduce a new paradigm of progressive multi-source domain adaptation (P-MSDA) for personalized facial expression recognition (FER), where the goal is to adapt to a single unlabeled target subject by progressively leveraging the most relevant source subjects. Unlike conventional UDA or MSDA methods that adapt at a generic domain level or aggregate all sources at once, our approach focuses on subject-level personalization, recognizing that the distribution of facial expression can vary significantly across individuals due to factors such as demographics, age, and cultural background. P-MSDA begins by ranking source subjects based on their similarity to the target in feature space and selecting only those above a threshold, ensuring that adaptation starts from the most relevant sources as illustrated in Fig. 1(b). As training progresses, we re-calculate similarities to dynamically incorporate new sources, thereby combining the principles of curriculum learning [23] (starting from “easier” sources) and self-paced learning (adapting as the model matures).
@@
To train such a model, a naive approach begins with the easiest source subject and gradually moves towards the hardest subjects while storing all the previously seen subjects. Another option is to train a model by ignoring the previously learned subjects and only training with the newly added source subject. The former approach requires significant computation power and memory consumption, while the latter setting will lean towards the problem of catastrophic forgetting [27, 28]. To avoid using every visited subject and eliminate the problem of forgetting, we follow the DIL replay-based [29] strategy, where we create a replay dictionary that preserves a small set of previously adapted samples based on the closest source points from the target clusters and only incorporates those with the newly added subject. We keep updating the replay dictionary as the training progresses. Our method aims to strike a balance between leveraging diverse source information and maintaining the domain-specific characteristics of the target.
@@
The main contributions of this paper are summarized as follows. (1) A novel progressive learning framework for MSDA personalized FER that exploits the closest source subjects for target adaptation. Additionally, we present an effective training strategy based on DIL that starts with the closest source and gradually introduces a new subject while limiting the number of source subjects to the top-N pertinent sources. This approach minimizes the risk of data corruption from mixed source domains and reduces computational complexity. (2) Inspired by the incremental learning replay technique, we present a density-based sample selection mechanism that retains the most relevant samples from previously visited source subjects, avoiding catastrophic forgetting. (3) The performance of the proposed method is extensively evaluated on diverse groups of source and target subjects using two benchmark pain estimation datasets, BioVid and UNBC-McMaster, as well as two in-the-wild affective computing datasets, Aff-Wild2 and Behavioral Ambivalence/Hesitancy (BAH). We also show the efficacy of our method by evaluating cross-dataset, specifically UNBC-McMaster (source) to BioVid (target). Further, we present a comprehensive analysis of the selection of previous samples for better target adaptation.
@@ II-B Domain Adaptation
Domain adaptation (DA) methods adapt from labeled source domains to unlabeled target domains, and can be broadly grouped into discrepancy-based [15, 13], reconstruction-based [30, 31], and adversarial-based [32, 33, 34] approaches. While effective, most DA methods rely on a single source domain, limiting their ability to generalize across diverse data distributions. To overcome this, our prior work on subject-based MSDA [18] introduced multiple source domains (subjects), showing improved performance by leveraging inter-subject diversity.
@@ II-C Multi-Source Domain Adaptation
Multi-source domain adaptation (MSDA) leverages multiple labeled source datasets to improve target domain accuracy and robustness. Existing MSDA approaches in image classification [19, 35, 36, 16] include adversarial [16, 36], self-supervised [37, 38], and discrepancy-based [19] strategies, which mitigate domain shift effectively on standard benchmarks [35]. However, their application to subject-based facial expression recognition (FER) remains limited, especially when scaling to many source domains. Zeeshan et al. [18] addressed this by adapting to a target subject from up to 77 sources, selecting the top-$k$ most similar sources for joint training. While effective, aggregating multiple sources at once can introduce noise from less relevant domains. In contrast, our method ranks sources by distributional similarity to the target and incorporates them progressively, allowing fine-grained adaptation while mitigating negative transfer from less relevant subjects.
@@ II-E Incremental Learning
Incremental learning (IL) sequentially acquires knowledge from multiple datasets without retaining full access to past data and is generally categorized into task-, class-, and domain-incremental learning (DIL)[43]. This work focuses on DIL, which adapts to new domains while avoiding catastrophic forgetting. DIL methods have been applied in image classification[25], autonomous driving [44], and semantic segmentation [45], and typically fall into three categories: parameter isolation [46], regularization-based approaches (e.g., knowledge distillation [47, 48]), and replay-based strategies [29, 28, 49]. For domain adaptation, prior works [50, 51] address multiple target domains incrementally, but often overlook domain shift. In contrast, our method adapts to a single unlabeled target by progressively introducing source domains while selectively preserving only the most relevant samples, thereby mitigating domain discrepancies and improving target personalization.
@@ II-F Self-paced Curriculum Learning
Curriculum learning (CL) [23] and self-paced learning (SPL) [24] start with easier tasks or samples and progressively move to more complex ones. CL leverages prior knowledge to guide the sequence, while SPL adjusts dynamically to the learner’s pace. Several works adopt these ideas for UDA, e.g., Choi et al. [52] select high-density target samples first, and Wang et al. [53] choose target samples agreed upon by multiple source classifiers. Jiang et al. [54] combine both paradigms into a unified self-paced curriculum learning (SPCL) framework. Yang et al. [55] apply CL to MSDA by selecting source samples similar to the target distribution. However, these approaches typically focus on confident target samples, without considering how source samples are introduced. More recently, Zhu et al. [56] introduced Tensorial Multiview Low-Rank High-Order Graph Learning (MLRGL), which progressively refines pseudo-labels via masked–unmasked consistency and propagates multiview features through a high-order graph. While conceptually related to our progressive pseudo-labeling, MLRGL targets generic multiview UDA, whereas our method addresses a problem of personalized multi-source adaptation where both subject similarity and progressive inclusion of source subjects are crucial.
Figure 2: Overview of our proposed progressive MSDA method for the adaptation to the target subject. Source Selection Phase: We estimate the similarity matrix between every source and target embedding, followed by ranking the sources from most to least similar subjects. Progressive Domain Adaptation Phase: Ranked sources are progressively incorporated through iterative training steps (Train Step-1, Train Step-2, …, Train Step-n). At each step, a new source subject is introduced and aligned with the target by calculating discrepancy and supervised losses. The Augmented Confident Pseudo-label (ACPL) technique from [18] generates reliable pseudo-labels for the target. Finally, we create a replay dictionary using a density-based selection to preserve previously visited relevant source samples.
@@ III-A Preliminaries
In MSDA, given a set of labeled source domains $\mathcal{S}={\{\mathbf{S}_{1},\mathbf{S}_{2},\ldots,\mathbf{S}_{a},\ldots,\mathbf{S}_{D}\}}$ and a single unlabeled target domain $\mathbf{T}$, where $a={\{1,2,\ldots,D}\}$ is the number of source domains. To preserve relevant samples from previous source domains, we define a replay domain $\mathbf{R}$. We define a source domain as $\mathbf{S}_{a}={\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})}\}_{i=1}^{N^{\text{s}}}$, where $N^{\text{s}}$ represents the number of samples within each source domain $a$. Assuming that $\mathbf{x}_{i}^{\text{s}}$ represents an input embedding of a source domain sample $i$ produced by an encoder $F$, the labeled source domain input space is defined as $\mathbf{x}_{i}^{\text{s}}$, and their respective labels as $y_{i}^{\text{s}}$. We define a single unlabeled target domain as $\mathbf{T}={\{\mathbf{x}_{i}^{\text{t}}\}}_{i=1}^{N^{\text{t}}}$, where $N^{\text{t}}$ denotes the number of samples within the target domain. A replay domain is defined as $\mathbf{R}={\{(\mathbf{x}_{i}^{\text{r}},y_{i}^{\text{r}})}\}_{i=1}^{N^{\text{r}}}$, where $N^{\text{r}}$ represents the number of relevant source domains from $\mathcal{S}$, which keeps on updating after adapting the model $\Theta$ to the target domain $\mathbf{T}$. The model $\Theta$ integrates the representation learning module $F$ with the classifier $C$. The objective is to leverage the information of source domains by gradually introducing domains from $\mathcal{S}$ to improve the performance of a model $\Theta$ on the target domain $\mathbf{T}$. Therefore, at each training step, the model $\Theta$ aims to learn from a new source domain from $\mathcal{S}$ and $\mathbf{R}$, which retain the most representative information from the previous source domains.